OpenStudy (anonymous):

If sin θ= 1/4, 0<θ<π/2, find the exact value of the following: sin(2θ), cos(2θ), sin(θ/2), cos(θ/2).

OpenStudy (anonymous):

\[\sin2\theta = 2\sin \theta.\cos \theta\] \[\cos2\theta=1-2\sin^2 \theta\] \[\sin ^2\frac{ \theta }{ 2 }=\frac{ 1-\cos \theta }{ 2 }\] \[\cos^2\frac{ \theta }{ 2 }\] first find cos theta value and plug these values in formulas

OpenStudy (anonymous):